(9x-54)(8x+32)=0 can be solved by using what property?
Answer by Flannery(124)

Do you need it in quadratic form or do you need solutions for X? You can FOIL it into quadratic form. Multiply the first two, outer two, inner two and last two terms together and add the inner two of the four resulting terms together to get 72x^2-144x-1728=0. Now you are in quadratic form. But to solve for values of x, set each bracketed term to equal zero and solve each for x.
9x-54=0;8x+32=0
by solving these two terms for x, you will have gotten (-4,6)
9x=54
x=6
and
8x=-32
x=-4